[Resolved] XN10 Packet loss

Starting at approximately 22:20 GMT+1, xn10 was experiencing high packet loss. Due to limited access to the server we gracefully shutdown all VM’s and brought them back up to make some configuration changes. Apologies for the reboot.

We have now identified one VM is the destination of a low bandwidth, high concurrency DoS attack which has now been null routed and we continue to monitor.

[Resolved] XN1 Outage

We are aware of an issue with our XN1 node, we have identified the issue and are working to resolve it as quickly as possible.

Updates will be provided here.

Update @ 16:52: DC Staff are running slow, frustrating but nothing we can do unfortunately to speed this up.

Update: This is now resolved.

[Resolved] xn3 outage

xn3.pcsmarthosting.co.uk currently has issues with the RAID controller which we are working to resolve. Further updates will follow.

Update: We’ve traced this to a bug in the upstream Linux 4.9 kernel. The raid controller and array appear to be healthy. We’ll bring the server back up on a secondary kernel to restore service. Further investigations will be carried out in our test environment.

[Resolved] XN17 – Planned Maintenance 2017/01/27 – 22:00GMT

Update @ 21:55 – This has now begun.

Update @ 22:28 – This has now been completed and VPS are now booting back up

This is an advanced notification of upcoming maintenance on hostnode XN17, please see below for further details.

When Is this happening?
It will begin on Friday, 27th January 2017 at 22:00 GMT

How long will it take?
We are hoping this should take no longer than 60 minutes.

Is any Downtime expected?
Yes, the server will need to be powered down so downtime is to be expected for the duration.

Updates will be posted here.

[Resolved] XN10 Outage

We have been made aware of an issue with XN10 by monitoring and are currently investigating the issue, further updates will be provided here when we have them

Update: This has now been resolved.

[Resolved] Web1 Read Only

Our web1 server has currently gone into a read only state and our technicians are currently investigating this.

Updates will be provided here when they’re available.

Update: Onsite staff are currently hooking up a crash cart to this machine.

Update: It appears a hard disk in the RAID10 array has failed, and caused the controller to hang. This is rare but it can happen. There are some filesystem inconsistencies we are working to repair then the server will be brought back online.

Update: Damage has been repaired and didn’t look too bad, just doing a 2nd pass now to be absolutely sure and will then boot.

Update: Server is now back online and the RAID array is rebuilding.

[Resolved] VZ1 Outage

We are currently aware of an issue with our VZ1 Node, we are investigating this and hope to have service restored in the next 30 minutes.

Update: This has now been resolved and the machine is back online.

[Resolved] Billing and Ticketing DDoS Outage

We are currently facing a large DDoS attack on the offsite DataCentre we use for our billing and ticketing systems, the incident is being treated as a priority 1 and we will post updates here as we get them.

We offload our billing and ticketing systems to a different DataCentre for redundancy so should our primary DataCentre be unavailable access to your billing and ticketing systems would be unaffected.

Update: This has now been mitigated and the issue resolved.

Because Uptime Matters